shift summary for DAY 7/23/99

Arun Tripathi (
Fri, 23 Jul 1999 13:29:47 -0700 (PDT)

Shift Summary for DAY shift, 23 July 99

On Shift: U. Erkarslan, A. Tripathi

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
No power to TeV.

Runs ended this shift:
Tapes brought to FCC this shift: None

Accesses during shift: None

Overview of shift:

No beam, TeV down upon ComEd orders.

Description of Shift:

Arun tried to find the reason for damp summary core dump. After
doing 10 inits under various conditions (killing the run in the
middle and doing DRT_FRESH_START etc.), one core dump was obtained.
It happened because dampfwdr (damp forwarder) process had died for
some unkown reason, and damp summary was trying to fetch info from
it. When a core dump does happen, one has to logout
of the Ktev3 window and then log back in to be able to initialize
a run successfully.

Cleaned all the tape drives.

Since no beam is expected before ~6, Sydeny is using these drives
for copying the wrongly labelled tapes.

SDAQ processes "vacuum" and "trd_gas" kept dying, even after a
fresh start of the sdaq. Theo called and said to reboot the
sdaq monitor right above the ktrig monitor. That seems to have
fixed it.

Elliott replaced the Azlat in crate 53, slot 13. All bits now work,
and no FFFF errors appear during run start.

MCR called to say that TeV startup will begin at 5pm, and
optimistically, beam to ktev can arrive 1-2 hrs later.

Follow up Notes on Previous Shift

Page Leonard Nelson (I guess day-time only?) if going on access.
He wants someone to check the humidity monitor.
Items to Follow up on Next Shift